### v1

# Fixing non-blocking data structures for better compatibility with memory reclamation schemes

Abstract
We present a new technique, Safe Concurrent Optimistic Traversals (SCOT), to address a well-known problem with Hazard Pointers (HP) related to optimistic traversals. Unlike Epoch-Based Reclamation (EBR), HP enables strong robustness properties but lacks support for some data structures such as Harris’ original list and Natarajan-Mittal tree, which can slow down operations substantially (e.g., when replacing with Harris-Michael list) or make correct implementations impossible altogether (e.g., Natarajan-Mittal tree).
Recent works propose replacing HP with another robust memory reclamation schemes that support optimistic traversals such as HP++. However, HP++ is generally slower than HP, has undesirable applicability trade-offs, and more complex implementation, among other issues. We propose a different method which keeps HP intact but instead relies on data structure adaptations. Moreover, our method applies to other recent (and faster) robust reclamation schemes such as Hazard Eras (HE) and Interval-Based Reclamation (IBR), which are HP-like and have similar problems related to optimistic traversals.
We implement two fundamentally different data structures: Harris’ list and Natarajan-Mittal tree. Our method enables their first correct implementations with optimistic traversals with HP and IBR. Our method is also directly applicable to other data structures which extend Harris’ list, e.g., skip lists and hash maps. We observe similar performance benefits of Harris’ original list (vs. Harris-Michael list) that were previously only available to EBR users. As we discuss in the paper, our Natarajan-Mittal tree is the first bug-free implementation for HP and IBR; it achieves very high throughput, comparable to that of EBR.
Keywords: hazard pointers, non-blocking, Harris’ list, Natarajan-Mittal tree
1 Introduction
With the growing prevalence of multicore and manycore systems, parallelization is becoming crucial for enhancing performance across the entire software stack – applications, middleware, core system services, and low-level components inside an operating system (OS). Non-blocking data structures, which are not dependent on mutual exclusion, are becoming increasingly popular due to their scalability, throughput, and latency characteristics. Unfortunately, many data structures that utilize non-blocking techniques conflict with fundamental assumptions of classical memory management. This is because memory cannot be promptly reclaimed after an object is removed from a data structure, as there may still exist stale pointers to the object due to ongoing operations.
When using manual memory management, memory has to be safely reclaimed by freeing memory only when all ongoing operations with stale pointers complete. A number of past researchers proposed various safe memory reclamation (SMR) techniques in C, C++, or Rust. Non-blocking SMR progress is often crucial for the overall progress guarantees, as we do not want to compromise data structure’s progress properties. Unfortunately, many SMR schemes are trying to attain several highly desirable properties simultaneously while often unknowingly compromising non-blocking progress or other important properties one way or the other. A recent work [21] sheds more light onto this conundrum: it proves that it is impossible to achieve three highly desirable SMR properties simultaneously, i.e., only two properties can ever be achieved: (A) robustness (bounded memory usage), (B) easy integration (no dependency on any special APIs, such as roll-back mechanisms or OS-based primitives), and (C) wide applicability (compatibility with most non-blocking data structures).
While (A) is non-negotiable for non-blocking progress, [^1] If memory usage is unbounded, no further system progress can be made once memory is exhausted, regardless of whether actual locks are present. past researchers would compromise either (B), (C), or other properties that are tangentially related to non-blocking progress, sometimes without fully realizing that. To make matters worse, many researchers prioritize (C) because, intuitively, their goal is to support as many data structures as possible. However, we argue that sacrificing (C) rather than (B) by carefully adapting existing problematic data structures is a “necessary evil”, which is a crucial new insight. First, (C) in its strongest form – applying to all data structures – only currently known to exist in blocking SMR schemes [21], which is a non-starter for us. Second, even if we accept (C) in its relaxed form – applying to most data structures – the absence of (B) would lead to either inability to achieve non-blocking progress in practice due to dependence on OS-based primitives, lack of universally-defined APIs, or both. (We discuss some specific examples in Section 5.) Third, the absence of (C) simply implies that adaptations are required, not impossibility of the implementation.
EBR [9] is a fast and easy-to-use approach but has a serious drawback that any stalled thread causes unbounded memory usage, i.e., it lacks (A). Once exhausting memory, the system must either reboot or block until memory is freed. Many researchers prefer EBR’s simplicity but attempt to solve its robustness issue. Unfortunately, such approaches typically sacrifice strict non-blocking progress [5, 22, 4, 1, 2], as discussed in Section 5, or need non-trivial roll-back mechanisms [20], which are fairly hard to generalize in terms of the API within the same library framework.
HP [14] is an antipode of EBR: it is somewhat slower and more difficult to use but provides robustness with strict non-blocking guarantees by placing a hard limit on the number of local pointers within each thread. Prior to this work, well-known data structures such as Harris’ linked list [10] with optimistic traversals [^2] The original paper [10] did not explicitly mention that search can be done in a read-only manner (w/o any CAS operations). In this paper, by Harris’ list, we denote the enhanced version [11] with optimistic read-only traversals. and Natarajan-Mittal’s binary search tree (BST) [15] were not known to work [^3] Despite presenting Natarajan-Mittal Tree results in [25, 17] for HP and other SMR schemes, it was later argued [3] that such implementations are buggy and leak memory. Moreover, [3] pointed out two bugs, and only one bug was resolved in [3]. Prior to our work, the second bug related to optimistic traversals remained unresolved. with HP [12, 3], which is due to HP’s lack of the wide-applicability (C) property. A number of other recent schemes, including Interval Based Reclamation (IBR) and Hazard Eras (HE), sometimes offer better performance but have similar issues to HP. HP++ [12] is another recent scheme. HP++ is slower than HP and, strictly speaking, makes compromises with both (B) and (C). However, HP++ still demonstrates support for these two data structures. We are inspired by HP++’s success with these data structures, which is indicative that many existing data structures can still be implemented, with an additional effort, despite prior beliefs regarding limited applicability of the SMR schemes in this category.
In this paper, we depart from a typical strategy of designing a “silver-bullet” SMR approach for the reasons described above. Instead, we argue that the existing problematic data structures can be modified to accommodate SMR schemes with strict non-blocking properties. We demonstrate our new method in Section 4 for Harris’ list, Natarajan-Mittal tree, and other data structures. In Section 5, we also touch upon how limitations of our approach with respect to wait-free traversals can be addressed.
2 Background
2.1 Lock- and Wait-Freedom
The most widely used non-blocking progress is lock-freedom, which allows threads to interfere with each other while still guaranteeing that at least one thread makes progress in a finite number of steps. This should not be confused with “lockless” approaches that simply avoid explicit locks, where one preempted thread can block all other threads. Wait-freedom is a much stronger property which requires that all threads make progress in a finite number of steps.
2.2 Memory Reclamation API
For both EBR and HP, the API is called directly. Each thread maintains a reservation, which is a globally accessible per-thread state. Memory objects, containing the required SMR header, are allocated using standard OS methods. Once a memory object is part of a data structure, it can only be reclaimed through two steps. First, after removing a pointer from the data structure, a memory object is retired. When the object becomes unreachable by any other concurrent thread, it is then returned to the OS.
EBR requires specifying both the start and end points of an operation to define the scope of protected memory accesses. In contrast, HP offers finer granularity in terms of reserved objects, requiring to explicitly wrap accesses to shared-memory objects. Additionally, in HP, the number of local pointers that can be held simultaneously by each thread is limited, and local pointers must be differentiated by unique indices.
2.3 Overview of Data Structures
In Table 1, we characterize common data structures based on their speed, (original) compatibility with SMR schemes, and compatibility when using our method. We omit hash maps as they are mostly arrays of Harris’ or Harris-Michael lists.
Table 1: Compatibility and performance characteristics of common non-blocking data structures.
Data Structure Performance EBR Original HP/IBR/HE HP/IBR/HE with SCOT
Harris’ Linked List [10]
Harris-Michael Linked List [13]
Fraser’s Skip List [9]
Herlihy-Shavit Skip List [11]
Natarajan-Mittal Tree [15]
Ellen et al. Tree [8]
These data structures employ “logical” deletion, i.e., a node inside a data structures is first simply marked (by stealing one bit from a node pointer), and then unlinked from the data structures (i.e., “physically” deleted). The original Harris’ List and Fraser’s Skip List use optimistic traversals, i.e., a chain of consecutive logically deleted nodes is deleted using one CAS operation; they allow implementing read-only search operations, which simply skip logically deleted nodes when iterating through the data structure. These optimizations improve performance and reduce contention.
However, as we discuss below, this approach is incompatible with HP. Harris-Michael List and Herlihy-Shavit Skip Lists use a different approach: they still use “logical” deletion. However, they mandate that logically deleted nodes are removed immediately after the first encounter and before proceeding further, including the search operation.
Ellen et al. Tree and Natarajan-Mittal Tree are more complex algorithms and typically achieve a higher overall throughput. There is a similar difference between them with respect to optimistic traversals. However, Natarajan-Mittal Tree’s ability to remove multiple logically deleted tree branches at the same time is even more crucial here – there is a higher chance of contention when multiple threads access the same node due to the hierarchical (rather than linear) structure of trees.
As we show in the table, data structures with optimistic traversals are incompatible with HP, IBR, and HE, a problem that we address in this paper.
2.4 Harris’ and Harris-Michael Linked Lists
The main challenge in lock-free linked lists is allowing a deleting thread to mark a node as deleted and update the preceding node’s next pointer. Timothy L. Harris solved this [10] by using helper threads during deletion. A deleting thread first marks a node as “logically” deleted by updating its next pointer, and then the node is “physically” removed by the deleting thread or a helper (conflicting concurrent operation). Harris’ list supports optimistic traversals, meaning that logically deleted nodes can still be traversed and bypassed by a search operation without immediate physical removal. This allows threads to defer unlinking logically deleted nodes until reaching the intended destination node.
However, this approach presents fundamental challenges for HP. Figure 1 shows an example with several concurrent operations. Suppose Thread 3 (not shown in the figure) has previously marked N2 for deletion but has not yet unlinked it. Then, Thread 2 marks N3 for deletion and unlinks the entire chain of nodes between N1 and N4 (assuming all consecutive nodes in the chain are logically deleted). Meanwhile, Thread 1 is traversing the list; it reaches N2 before Thread 2 physically removes the chain. Although N2 is now retired by Thread 2, it still remains in the HP limbo list. The reason for that is that Thread 1 has made a reservation prior to its physical deletion. N3 is also retired by Thread 2, but because Thread 1 never reserves N3, it is not guaranteed to stay in the HP limbo list, i.e., N3 is returned to the OS. Consequently, Thread 1 fails when accessing N3.
Maged M. Michael modified [13] Harris’ approach to resolve compatibility issues with HP: when a thread encounters a node marked as logically deleted during traversal, it immediately attempts to physically remove the node from the list.
Why does Michael’s approach work? Recall that accessing N2 is still safe, because its HP reservation is made prior to N2’s physical deletion. (To successfully make a reservation, HP verifies that the pointer from N1 to N2 remains intact.) Once N2 is physically deleted, it is retired and ends up in the HP limbo list. If we now attempt to reserve N3, HP falsely succeeds as the pointer from N2 to N3 remains intact unlike the pointer from N1 to N2. Thus, Michael’s approach is to ensure that pointers always change during physical removal, which happens automatically when we remove one node at a time. [^4] Another way to think about Michael’s approach is that if we remove nodes one by one, N3 takes place of N2, i.e., right after N1. N3 will be the first logically deleted node in the chain when making the HP reservation, which is still safe to access.
While it fully solves the HP problem, Michael’s approach increases the number of CAS operations and may lead to a higher contention among threads. It also makes read-only traversals impossible.
Figure 1: Unsafe optimistic traversal of Harris’ linked list with HP: accessing the first logically deleted node (N2) is safe but subsequent nodes may cause SEGFAULTs.
1 struct {
    node_t * Next;
     // Next node pointer
    key_t Key;
     // Any key type
2
3    } node_t;
   node_t * Head;
     // Head of the list
4    void Init()
5        node=malloc(sizeof(node_t));
6        node->Next = nullptr;
7        Head = node;
8
9   bool Insert(key_t key)
10        new = malloc(sizeof(node_t));
11        new->Key = key;
12        node_t **prev, *curr, *next;
13        while true do
14            if ( Do_Find(key, &prev, &curr, &next, false) )
15                free(new);
16                return false;
17
18           new->Next = curr;
19            if ( CAS(prev, curr, new )  return true;
20
21
22   bool Delete(key_t key)
23        node_t *curr, *next;
24        while true do
25            if ( !Do_Find(key, &prev, &curr, &next, false) )
26                return false;
27
28           if ( !CAS(&curr->next, next, getMarked(next) )  continue;
29            if ( CAS(prev, curr, next )  smr_retire(curr);
30            return true;
31
32
28 bool Do_Find(key_t key, node_t *** p_prev, node_t ** p_curr, node_t ** p_next, bool is_search)
29    node_t **prev, *curr, *next;
30    prev_next = nullptr;
31    prev = &Head;
32    curr = Head;
33    while true do
34        if ( curr == nullptr )  break ;
35        next = curr->Next;
36        if ( !isMarked(next) )
37            if ( curr->Key $\geq$ key )  break ;
38            prev = &curr->Next;
39            prev_next = next;
40
41       curr = getUnmarked(next);
42
43   if ( !is_search && prev_next != curr ) // Skip for Search
44        if ( !CAS(&prev, prev_next, curr )  goto 2 ;
45        do // Retire nodes from the deleted chain
46            node_t *node = getUnmarked(prev_next->next);
47            smr_retire(prev_next);
48            prev_next = node;
49
50       while (prev_next != curr);
51   *p_curr = curr;
52    *p_prev = prev;
53    *p_next = next;
54    return curr && (curr->Key == key);
55
56bool Search(key_t key)
57    node_t **prev, *curr, *next;
58    return Do_Find(key, &prev, &curr, &next, true);
59
Figure 2: Harris’ Linked List with optimistic traversals (ignoring memory reclamation).
In Figure 2, we demonstrate Harris’ list without the HP reclamation. The Insert method inserts a new node into the list. Before inserting a new node, Insert checks if the key of the to-be-inserted node exists via Do_Find. The new node gets inserted unless its key is already present in the list. Delete removes a node from the list. If the key is found in the list, the node is logically deleted at L25 by marking next, and then one attempt is made to unlink it from the list at L26.
In Search, the underlying Do_Find operation skips L41. In Insert and Delete, we also need to locate the right position for the key. Moreover, Insert and Delete ensure that we are not modifying pointers nodes which belong to nodes which are already (logically) deleted by cleaning up the discovered chain of logically deleted nodes in L41-L47 first. If we simply integrate HP without any modifications, L44 may crash while retrieving the next pointer in a loop.
3 Natarajan-Mittal Tree
Natarajan-Mittal Tree [15] is a fairly complex algorithm and its detailed discussion and pseudocode is beyond the scope of this paper; we describe its high-level idea, which suffices to explain how to apply SCOT to this tree. The tree keeps left and right pointers in every node. The actual keys and values are stored in leaf nodes, while internal nodes are simply guiding in which direction (left or right) the underlying search operation has to descend to the next level. A concept similar to logical deletion also exists in this tree. There are two types of marking: “flagging” when marking the leaf node for its logical deletion followed by “tagging” of its sibling node.
When performing the underlying search operation, we keep the successor node which is the last untagged node as well as its preceding node – ancestor. We also keep the immediate parent of the leaf node. The algorithm makes a very crucial observation: a chain of consecutively tagged edges can be eliminated with one CAS operation by updating ancestor’s link from successor to the remaining leaf node as shown in Figure 3.
Figure 3: Natarajan-Mittal tree: entire tagged edges can be removed.
4 Design
4.1 Bird’s-Eye View
The crux of the problem with the HP implementation is its inability to properly track physically deleted nodes in the original Harris’ list (or tagged edges in Natarajan-Mittal tree) while traversing through logically deleted nodes. However, if there was some way to confirm that the following logically deleted node is not yet physically deleted, optimistic traversals are still feasible.
In Figure 1, N2 is still safe to access because it is not physically deleted, i.e., N1 still points to N2. Moreover, according to Figure 2, we always make sure that the node to the left is not logically deleted and thus never unlink nodes in the middle of the chain, i.e., we can delete just N2, or both N2-N3, but we never delete N3 while keeping N2 in the list. In other words, we either remove the entire chain of logically deleted nodes or a subchain with the truncated tail.
Combining these observations together, we make one crucial insight that helps solve the issue with HP: It is still safe to access N3 and the following nodes as long as at every step, i.e., after making an HP reservation for the next logically deleted node, we verify that N1 still points to N2. In other words, once we reach N2, we declare it to be a “dangerous zone”. Until we reach the end of the chain of consecutive logically deleted nodes, we are going to stay in the dangerous zone and perform additional checks. More specifically, after retrieving N3’s pointer and creating its HP reservation, we check that N1 still points to N2. We will continue to perform these checks at every iteration until we reach the end of the chain. If the check fails, we cannot proceed further and need to restart the search operation from the very beginning.
4.2 SCOT for Harris’ List
In Figure 4, we present our approach for Harris’ list. Compared to Michael’s approach, we need an additional hazard pointer index (Hp3) which protects the first unsafe, i.e., logically deleted, node. This extra hazard pointer prevents the ABA problem, i.e., a case when the unsafe node gets recycled while we are traversing the list. By holding this extra hazard pointer, we can solely rely on pointer comparison. In other words, Hp3 protects N2 from our example above irrespective of where we currently are in the chain of logically deleted nodes. Like in Michael’s approach, we need Hp2, which protects the prev area. We note that in Harris’ list, it will be the last safe, i.e., not logically deleted, node.
The HP reclamation provides protect to safely retrieve the object and make a corresponding reservation. Additionally, we implement duplicate to duplicate the existing hazard pointer when moving to the next iteration. It is crucial to duplicate hazard pointers such that the old HP index has a lower numerical value than the new HP index (e.g., Hp0 to Hp1). This allows to avoid a small race window when iterating in the same (ascending) order of indices in retire.
Once we get to the dangerous zone of logically deleted nodes, Figure 4 runs a check in L16 which will ensure that reservation made in L8 is correct. Once the chain (or its subchain) gets unlinked, L16 fails and we restart the operation from the very beginning. We highlight in yellow color all major changes related to SCOT in HP.
1 bool Do_Find(key_t key, node_t *** p_prev, node_t ** p_curr, node_t ** p_next, bool is_search)
2    node_t **prev, *curr, *next;
3    prev_next = nullptr;
4    prev = &Head;
5    curr = hp.protect(Head, Hp1);
6    while true do
7        if ( curr == nullptr )  break ;
8        next = hp.protect(curr->Next, Hp0);
9        if ( !isMarked(next) )
10            if ( curr->Key $\geq$ key )  break ;
11            prev = &curr->Next;
12            prev_next = next;
            hp.duplicate(Hp1, Hp2);
             // Duplicate Hp1 (curr) to Hp2
              hp.duplicate(Hp0, Hp3);
             // Duplicate Hp0 (next) to Hp3
13
14       else
            // Dangerous zone (marked): Validate that last safe node (N1) still points to the first unsafe (N2) node
15              if ( *prev != prev_next ) goto 4;
16
17       curr = getUnmarked(next);
        hp.duplicate(Hp0, Hp1);
         // Duplicate Hp0 (next) to Hp1
18
19
// <—— Continuation of Do_Find() ——>
19
20    if ( !is_search && prev_next != curr ) // Skip for Search
21        if ( !CAS(&prev, prev_next, curr )  goto …;
22        ;
23        do // Retire nodes from the deleted chain
24            node_t *node = getUnmarked(prev_next->next);
25            smr_retire(prev_next);
26            prev_next = node;
27
28       while (prev_next != curr);
29   *p_curr = curr;
30    *p_prev = prev;
31    *p_next = next;
32    return curr && (curr->Key == key);
33
34
// Hazard Pointer Indices
const int Hp0 = 0;
  // Protects next
const int Hp1 = 1;
  // Protects curr
const int Hp2 = 2;
  // Protects last safe node (prev)
  const int Hp3 = 3;
  // Protects first unsafe node
Figure 4: SCOT for Harris’ list using HP reclamation: showing a new version of Do_Find only.
Figure 5: SCOT for Harris’ list: validating *prev = prev_next at every iteration while traversing through the dangerous zone.
4.3 SCOT for Natarajan-Mittal Tree
For the tree, we allocate 5 hazard pointers to protect nodes in the underlying search procedure: current, leaf, parent, successor, and ancestor. The current node points to the lowest node that is currently considered.
To properly support optimistic traversals, after each HP reservation of the current node, if the corresponding node is flagged or tagged, we need verify that ancestor still points to successor. If ancestor points to some other node or successor becomes tagged, we need to restart from the very beginning.
This idea is largely similar to the above-mentioned Harris’ list approach, except that no additional hazard pointer is required, as we already need to protect both ancestor and successor nodes. Additionally, we need to check both tagged and flagged conditions for the current node, as (flagged) leaf nodes’s hazard pointer reservations need to be also verified.
Due to complexity of the tree algorithm and very similar checks for optimistic traversals, we omit full pseudocode changes in this paper.
4.4 SCOT for Hash Map and Skip Lists
SCOT for these data structures is largely identical to Section 4.2. Hash Maps are directly based on linked lists. Fraser’s skip list uses the same idea of logically deleted nodes as Harris’ list.
5 Related Work
5.1 Non-Blocking SMR
The ERA theorem [21] demonstrates that at most two out of three properties – (A) robustness, (B) easy integration, and (C) wide applicability – can be achieved simultaneously. Moreover, strong applicability, as defined by that work, currently seems to be available only with SMR schemes that implement some kind of quiescence, similar to EBR. However, quiescence periods inevitably lead to potentially unbounded memory usage (i.e., lack of non-blocking progress).
A number of schemes, e.g., DEBRA+ [5], QSense [4], ThreadScan [2], ForkScan [1], NBR [22], propose a solution to EBR’s robustness (A) problem. However, they all need special OS-based support such as interaction with the OS scheduler or signal delivery. These OS mechanisms inevitably use locks in most typical OSs. They satisfy (A) and (C) properties but not (B), which is essential to the strict non-blocking progress. Other schemes that satisfy (A) and (C), e.g., VBR [20], need non-trivial roll-back mechanisms, which are too complicated to generalize in terms of the uniform library API.
There were also recent attempts to build non-blocking data structures using automatic garbage collection techniques. Generally speaking, garbage collection is subject to the same trade-offs and limitations, discussed above. FreeAccess [6] addresses the problem with (C) and optimistic traversals but forgoes (B) by requiring to divide code into separate read and write phases. OrcGC [7] is another lock-free garbage collector. OrcGC only works with data structures without loops and requires mechanisms similar to smart pointers, either in the compiler or the language itself; it can also be slower than HP in certain tests.
Other SMR schemes such as interval-based reclamation (IBR) [25], hazard eras (HE) [19], wait-free eras (WFE) [16], Hyaline [17], and Crystalline [18] provide HP-like interface; they still lack the wide-applicability property (C) but are often faster than HP. The method proposed in this paper to mitigate the problem with (C) is suitable to all of them.
5.2 Non-Blocking Data Structures
Timothy L. Harris [10] was the first to present a practical non-blocking lock-free linked list. Maurice Herlihy and Nir Shavit pointed out [11] that read-only optimistic search operations are feasible with Harris’ list. Unfortunately, Harris’ list is incompatible with HP [7, 6, 12, 3, 18, 17]. Maged M. Michael proposed [13] to disable optimistic traversals by unlinking logically deleted nodes one by one, making it feasible to use the list with HP. Our work allows to support Harris’ original list with HP and take advantage of optimistic traversals.
A skip list is a data structure with an efficient search, insertion, and deletion operations, achieving (probabilistically) logarithmic time complexity. Lock-free versions of skip lists [11, 9] are very similar to Harris’ list and Harris-Michael list, except that they effectively consist of several sublists. Consequently, challenges related to logical deletion and optimistic traversals are identical.
Hash maps store and retrieve key-value pairs in constant time. A lock-free hash map [13] is, effectively, an array of linked lists, which can be based on either Harris’ or Harris-Michael approach.
Binary search trees are also commonly used for efficient search, insertion, and deletion operations. Ellen at al. Tree [8] implements one such approach which is compatible with HP. However, Natarajan-Mittal Tree [15] is significantly faster (around 43% according to [12]) due to its ability to remove multiple concurrently tagged edges in one pass. Natarajan-Mittal Tree is a very popular data structure and evaluated in several recent works [25, 3, 17, 12]. Prior to our work, as pointed out in [3, 12], Natarajan-Mittal Tree was incompatible with HP, IBR, and other similar schemes.
5.3 Wait-Free Optimistic Traversals
With our approach, optimistic traversals might have to occasionally restart the search operation from the beginning if, due to overlapping modifications, the state of the data structure diverges significantly and cannot be recovered locally. This guarantees lock-free but not wait-free progress. This limitation is very similar to that of HP++’s existing solution which also does not support wait-free optimistic traversals due to potential restarts.
However, this restart is inherently unavoidable with data structures that use optimistic traversals: it is still needed for insertion and deletion operations, regardless of the search operation or the reclamation scheme used. We note that we can still implement wait-free search traversals by using Timnat-Petrank’s method [23, 24] which allows restarts and works with HP. This method can be applied to various data structures, including linked lists. Timnat-Petrank’s original wait-free linked list, implemented in HP, is similar to Harris-Michael list. With our method, it can now leverage Harris’ optimistic traversals for faster performance.
6 Conclusion
We introduced safe concurrent optimistic traversals (SCOT), a new method which enables support for data structures that are normally incompatible with HP, IBR, HE, and other similar methods that lack the wide applicability support. Despite prior beliefs [12] of incompatibility of Natarajan-Mittal Tree and Harris’ List with HP, we have found that not only they are feasible with HP, IBR, and HE, but they also can largely retain performance benefits when comparing to equivalent data structures without optimistic traversals.
We hope that our work will help to spur further research and reevaluate the existing challenges related to safe memory reclamation because lack of the wide applicability property may be addressed by other means.
Availability
SCOT’s code and benchmark for Harris’ list and Natarajan-Mittal tree is available at https://github.com/rusnikola/scot

### v2

@@
 Abstract
-We present a new technique, Safe Concurrent Optimistic Traversals (SCOT), to address a well-known problem with Hazard Pointers (HP) related to optimistic traversals.
-Unlike Epoch-Based Reclamation (EBR), HP enables strong robustness properties but lacks support for some data structures such as Harris’ original list and Natarajan-Mittal tree, which can slow down operations substantially (e.g., when replacing with Harris-Michael list) or make correct implementations impossible altogether (e.g., Natarajan-Mittal tree).
-Recent works propose replacing HP with another robust memory reclamation schemes that support optimistic traversals such as HP++.
-However, HP++ is generally slower than HP, has undesirable applicability trade-offs, and more complex implementation, among other issues.
-We propose a different method which keeps HP intact but instead relies on data structure adaptations.
-Moreover, our method applies to other recent (and faster) robust reclamation schemes such as Hazard Eras (HE) and Interval-Based Reclamation (IBR), which are HP-like and have similar problems related to optimistic traversals.
-We implement two fundamentally different data structures: Harris’ list and Natarajan-Mittal tree.
-Our method enables their first correct implementations with optimistic traversals with HP and IBR.
-Our method is also directly applicable to other data structures which extend Harris’ list, e.g., skip lists and hash maps.
-We observe similar performance benefits of Harris’ original list (vs.
-Harris-Michael list) that were previously only available to EBR users.
-As we discuss in the paper, our Natarajan-Mittal tree is the first bug-free implementation for HP and IBR; it achieves very high throughput, comparable to that of EBR.
+We present a new technique, Safe Concurrent Optimistic Traversals (SCOT), to address a well-known problem related to optimistic traversals with both classical and more recent memory reclamation schemes, such as Hazard Pointers (HP), Hazard Eras (HE), Interval-Based Reclamation (IBR), and Hyaline.
+Unlike Epoch-Based Reclamation (EBR), these schemes guarantee protection against stalled threads (robustness) but lack support for well-known data structures with optimistic traversals such as Harris’ original list, Natarajan-Mittal tree, among others.
+For these reclamation schemes, existing data structure implementations are either buggy (e.g., Natarajan-Mittal tree) or come with performance trade-offs (e.g., Harris-Michael modified list).
+A recent work, HP++, supports optimistic traversals but uses a different API and is generally slower than even HP, not to mention more recent schemes such as IBR or Hyaline.
+Moreover, it has undesirable applicability trade-offs, and more complex implementation, among other issues.
+We propose a different method which keeps existing reclamation schemes intact but instead relies on data structure adaptations.
+Unlike existing Harris-Michael approach or HP++, our method retains performance benefits of the original data structure and also does not compromise performance of the underlying reclamation scheme.
+In fact, for IBR and Hyaline, our results almost match those of EBR, which often serves as a practical upper bound due to its great performance.
+We implement and evaluate two fundamentally different data structures: Harris’ list and Natarajan-Mittal tree.
+SCOT enables their first correct implementations with optimistic traversals for HP, HE, IBR, and Hyaline.
+Our method is also directly applicable to other data structures, e.g., Fraser’s skip list or hash map that uses Harris’ list internally.
+We observe similar performance speedups (e.g., Harris vs.
+Harris-Michael list) that were previously available only to EBR users.
+As we discuss in the paper, our version of Natarajan-Mittal tree is the first bug-free implementation for HP, HE, IBR, and Hyaline; it achieves very high throughput, comparable to that of EBR.
 Keywords: hazard pointers, non-blocking, Harris’ list, Natarajan-Mittal tree
@@
 A number of past researchers proposed various safe memory reclamation (SMR) techniques in C, C++, or Rust.
-Non-blocking SMR progress is often crucial for the overall progress guarantees, as we do not want to compromise data structure’s progress properties.
+Non-blocking SMR progress is often crucial for the overall progress guarantees, as we do not want to compromise data structure progress properties.
 Unfortunately, many SMR schemes are trying to attain several highly desirable properties simultaneously while often unknowingly compromising non-blocking progress or other important properties one way or the other.
 A recent work [22] sheds more light onto this conundrum: it proves that it is impossible to achieve three highly desirable SMR properties simultaneously, i.e., only two properties can ever be achieved: (A) robustness (bounded memory usage), (B) easy integration (no dependency on any special APIs, such as roll-back mechanisms or OS-based primitives), and (C) wide applicability (compatibility with most non-blocking data structures).
-While (A) is non-negotiable for non-blocking progress, [^1] If memory usage is unbounded, no further system progress can be made once memory is exhausted, regardless of whether actual locks are present.
+While striving to have (A) for non-blocking progress, [^1] If memory usage is unbounded, no further system progress can be made once memory is exhausted, regardless of whether actual locks are present.
 past researchers would compromise either (B), (C), or other properties that are tangentially related to non-blocking progress, sometimes without fully realizing that.
@@
 (We discuss some specific examples in Section 6.) Third, the absence of (C) simply implies that adaptations are required, not impossibility of the implementation.
-EBR [9] is a fast and easy-to-use approach but has a serious drawback that any stalled thread causes unbounded memory usage, i.e., it lacks (A).
+EBR [9] is a fast and easy-to-use approach but has a serious drawback that any stalled thread results in unbounded memory usage, i.e., it lacks (A).
 Once exhausting memory, the system must either reboot or block until memory is freed.
@@
 with HP [12, 3], which is due to HP’s lack of the wide-applicability (C) property.
-A number of other recent schemes, including Interval Based Reclamation (IBR) and Hazard Eras (HE), sometimes offer better performance but have similar issues to HP.
+A number of other recent schemes, including Interval Based Reclamation (IBR) [26], Hyaline-1S [18], and Hazard Eras (HE) [20], sometimes offer better performance but have similar issues to HP.
 HP++ [12] is another recent scheme.
@@
 We are inspired by HP++’s success with these data structures, which is indicative that many existing data structures can still be implemented, with an additional effort, despite prior beliefs regarding limited applicability of the SMR schemes in this category.
-In this paper, we depart from a typical strategy of designing a “silver-bullet” SMR approach for the reasons described above.
+This paper makes the following contributions.
+First, we depart from the typical strategy of designing a “silver-bullet” SMR approach for the reasons described above.
 Instead, we argue that the existing problematic data structures can be modified to accommodate SMR schemes with strict non-blocking properties.
 We demonstrate our new method in Section 3 for Harris’ list, Natarajan-Mittal tree, and other data structures.
-In Section 5, we also touch upon how limitations of our approach with respect to wait-free traversals can be addressed.
+In Section 5, we evaluate these data structures with HP, HE, IBR and Hyaline-1S.
+We also compare results to EBR and show that performance benefits can be preserved in the same manner.
+Finally, in Section 6, we also touch upon how limitations of our approach with respect to wait-free traversals can be addressed.
 2 Background
@@
 The most widely used non-blocking progress is lock-freedom, which allows threads to interfere with each other while still guaranteeing that at least one thread makes progress in a finite number of steps.
-This should not be confused with “lockless” approaches that simply avoid explicit locks, where one preempted thread can block all other threads.
+(It should be noted that good lock-free algorithms allow multiple threads to make concurrent progress in practice.) This should not be confused with “lockless” approaches that simply avoid explicit locks, where one preempted thread can block all other threads.
 Wait-freedom is a much stronger property which requires that all threads make progress in a finite number of steps.
@@
 In Table 1, we characterize common data structures based on their speed, (original) compatibility with SMR schemes, and compatibility when using our method.
-We omit hash maps as they are mostly arrays of Harris’ or Harris-Michael lists.
-Table 1: Compatibility and performance characteristics of common non-blocking data structures.
-Data Structure Performance EBR Original HP/IBR/HE HP/IBR/HE with SCOT
+We omit hash maps as they are simply arrays of Harris’ or Harris-Michael lists.
+Table 1: Subjective performance characterization of common non-blocking data structures and their compatibility with SMR schemes.
+Data Structure Fast EBR HP/HE/IBR/Hyaline-1S HP/HE/IBR/Hyaline-1S with SCOT
 Harris’ Linked List [10]
@@
 Ellen et al. Tree [8]
-These data structures employ “logical” deletion, i.e., a node inside a data structures is first simply marked (by stealing one bit from a node pointer), and then unlinked from the data structures (i.e., “physically” deleted).
+These data structures employ “logical” deletion, i.e., a node inside a data structure is first simply marked (by stealing one bit from a node pointer), and then unlinked from the data structures (i.e., “physically” deleted).
 The original Harris’ List and Fraser’s Skip List use optimistic traversals, i.e., a chain of consecutive logically deleted nodes is deleted using one CAS operation; they allow implementing read-only search operations, which simply skip logically deleted nodes when iterating through the data structure.
@@
 However, Natarajan-Mittal Tree’s ability to remove multiple logically deleted tree branches at the same time is even more crucial here – there is a higher chance of contention when multiple threads access the same node due to the hierarchical (rather than linear) structure of trees.
-As we show in the table, data structures with optimistic traversals are incompatible with HP, IBR, and HE, a problem that we address in this paper.
+As we show in the table, data structures with optimistic traversals are incompatible with HP, HE, IBR, and Hyaline-1S, a problem that we address in this paper.
 2.4 Harris’ and Harris-Michael Linked Lists
@@
 4    void Init()
-5        node=malloc(sizeof(node_t));
-6        node->Next = nullptr;
-7        Head = node;
+5        Head = malloc(sizeof(node_t));
+6        Head->Next = nullptr;
+7        Head->Key = $\infty$;
 8
@@
 22   bool Delete(key_t key)
-23        node_t *curr, *next;
+23        node_t **prev, *curr, *next;
 24        while true do
@@
 32
-28 bool Do_Find(key_t key, node_t *** p_prev, node_t ** p_curr, node_t ** p_next, bool is_search)
+28 bool Do_Find(key_t key, node_t ***p_prev, node_t **p_curr, node_t **p_next, bool srch)
 29    node_t **prev, *curr, *next;
@@
 42
-43   if ( !is_search && prev_next != curr ) // Skip for Search
+43   if ( !srch && prev_next != curr ) // Search: skip
 44        if ( !CAS(&prev, prev_next, curr )  goto 2 ;
@@
 46            node_t *node = getUnmarked(prev_next->next);
-47            smr_retire(prev_next);
+47            smr.retire(prev_next);
 48            prev_next = node;
@@
 In Search, the underlying Do_Find operation skips L41.
-In Insert and Delete, we also need to locate the right position for the key.
-Moreover, Insert and Delete ensure that we are not modifying pointers nodes which belong to nodes which are already (logically) deleted by cleaning up the discovered chain of logically deleted nodes in L41-L47 first.
-If we simply integrate HP without any modifications, L44 may crash while retrieving the next pointer in a loop.
-3 Natarajan-Mittal Tree
+In Insert and Delete, Do_Find locates the right position for the key.
+Moreover, Insert and Delete ensure that we are not modifying the pointers of nodes that have been (logically) deleted by cleaning up the identified chain of logically deleted nodes in L41-L47 first.
+If we simply integrate HP without any changes, L44 may crash while retrieving the next pointer in a loop.
+2.5 Natarajan-Mittal Tree
 Natarajan-Mittal Tree [16] is a fairly complex algorithm and its detailed discussion and pseudocode is beyond the scope of this paper; we describe its high-level idea, which suffices to explain how to apply SCOT to this tree.
@@
 We note that in Harris’ list, it will be the last safe, i.e., not logically deleted, node.
-The HP reclamation provides protect to safely retrieve the object and make a corresponding reservation.
-Additionally, we implement duplicate to duplicate the existing hazard pointer when moving to the next iteration.
+The HP reclamation provides a protect function to safely retrieve an object and make the corresponding reservation.
+Additionally, we implement dup to duplicate the existing hazard pointer when moving to the next iteration.
 It is crucial to duplicate hazard pointers such that the old HP index has a lower numerical value than the new HP index (e.g., Hp0 to Hp1).
 This allows to avoid a small race window when iterating in the same (ascending) order of indices in retire.
-Once we get to the dangerous zone of logically deleted nodes, Figure 4 runs a check in L16 which will ensure that reservation made in L8 is correct.
-Once the chain (or its subchain) gets unlinked, L16 fails and we restart the operation from the very beginning.
-We highlight in yellow color all major changes related to SCOT in HP.
-1 bool Do_Find(key_t key, node_t *** p_prev, node_t ** p_curr, node_t ** p_next, bool is_search)
+(An alternative to dup would be index renaming via an indirection array.
+From our empirical observations, we found duplication to be generally cheaper.)
+One tricky part is to avoid a premature duplication into Hp3, which would induce a memory barrier not present in Harris-Michael approach, potentially making the algorithm more costly.
+We achieve that by unrolling the loop in the original Harris’ algorithm and splitting it into two phases: (1) iteration through the safe zone and (2) iteration through the dangerous zone.
+Phase 1 only duplicates (shifts) prev and curr hazard pointers from curr and next, correspondingly.
+This is analogous to Harris-Michael algorithm.
+Upon leaving Phase 1, curr is duplicated (Hp3) in L18.
+After going to Phase 2, we no longer duplicate curr into prev (Hp2), which give us an extra benefit compared to Harris-Michael approach as the number of memory barriers through the dangerous zone is reduced.
+1 bool Do_Find(key_t key, node_t ***p_prev, node_t **p_curr, node_t **p_next, bool srch)
 2    node_t **prev, *curr, *next;
@@
 5    curr = hp.protect(Head, Hp1);
+6    next = hp.protect(curr->Next, Hp0);
 7    while true do
-7        if ( curr == nullptr )  break ;
+8        do
+9            if ( curr && curr->Key $\geq$ key )  goto 4 ;
+10            prev_next = nullptr;
+11            prev = &curr->Next;
+            hp.dup(Hp1, Hp2);
+             // Duplicate Hp1[curr] to Hp2
+12            curr = getUnmarked(next);
+13            if ( curr == nullptr )  goto 4 ;
+            hp.dup(Hp0, Hp1);
+             // Duplicate Hp0[next] to Hp1
 14            next = hp.protect(curr->Next, Hp0);
-9        if ( !isMarked(next) )
-10            if ( curr->Key $\geq$ key )  break ;
-11            prev = &curr->Next;
-12            prev_next = next;
-            hp.duplicate(Hp1, Hp2);
-             // Duplicate Hp1 (curr) to Hp2
-              hp.duplicate(Hp0, Hp3);
-             // Duplicate Hp0 (next) to Hp3
-13
-14       else
-            // Dangerous zone (marked): Validate that last safe node (N1) still points to the first unsafe (N2) node
+15
+16       while !isMarked(next);
+         prev_next = hp.dup(Hp1,Hp3);
+         // Hp1[curr] to Hp3
+        // Dangerous zone (marked): Check last safe node (N1) still points to the first unsafe (N2) node
+17          do
+18              curr = getUnmarked(next);
+19              if ( curr == nullptr )  goto 4 ;
+              hp.dup(Hp0, Hp1);
+             // Duplicate Hp0[next] to Hp1
+20              next = hp.protect(curr->Next, Hp0);
 21              if ( *prev != prev_next ) goto 4;
-16
-17       curr = getUnmarked(next);
-        hp.duplicate(Hp0, Hp1);
-         // Duplicate Hp0 (next) to Hp1
-18
-19
+22
+23       while isMarked(next);
+24
 // <—— Continuation of Do_Find() ——>
-19
-20    if ( !is_search && prev_next != curr ) // Skip for Search
-21        if ( !CAS(&prev, prev_next, curr )  goto …;
-22        ;
+26
+    // Skip for search
+27    if ( !srch && prev_next != nullptr && prev_next != curr )
+28        if ( !CAS(&prev, prev_next, curr )
+29            goto 4;
+30
 31       do // Retire nodes from the deleted chain
-24            node_t *node = getUnmarked(prev_next->next);
+32            node_t node;
+33            node = getUnmarked(prev_next->next);
 34            smr_retire(prev_next);
 35            prev_next = node;
-27
+36
 37       while (prev_next != curr);
@@
 41    return curr && (curr->Key == key);
-33
-34
+42
+43
 // Hazard Pointer Indices
@@
 Figure 4: SCOT for Harris’ list using HP reclamation: showing a new version of Do_Find only.
+Once we get to the dangerous zone of logically deleted nodes, Figure 4 runs a check in L24 which will ensure that reservation made in L23 is correct.
+Once the chain (or its subchain) gets unlinked, L24 fails and we restart the operation from the very beginning.
+We highlight in the green color all major changes related to SCOT in HP.
+We use pink to highlight the dangerous zone traversal.
+3.2.1 Recovery Optimization
+For simplicity, we have stated that when validation fails, we need to go back to the beginning of the list.
+While this is acceptable, this validation check is not always critical: the last safe node may simply point to another node now (e.g., when a new node was inserted or the chain of logically deleted nodes was already eliminated by a concurrent thread).
+In such cases, we can simply escape from the dangerous zone and continue to new node.
+We must, however, ensure that the last safe node is still not logically deleted.
+When it is deleted, the last safe node (prev) is in a dangerous zone itself (though still safe to access due to prior reservation).
+Depending on the SMR algorithm, we have different choices.
+For HP and HE, protect uses indices and cancels previous reservations.
+As keeping a predecessor of prev would induce extra hazard pointers and duplication (hence memory barriers), it makes sense to simply go back to the beginning of the list when prev gets logically deleted.
+For IBR and Hyaline-1S, protect has a cumulative effect and does not cancel previous reservations.
+For these schemes, we implement a small on-stack ring buffer of prev pointers, which allows us to fall further back in the list when prev gets logically deleted.
+Empirically, we found that a ring buffer of size 8 is more or less optimal.
+Only when exhausting all 8 predecessors, do we go back to the beginning of the list.
+(See Figure 6.)
 Figure 5: SCOT for Harris’ list: validating *prev = prev_next at every iteration while traversing through the dangerous zone.
+Figure 6: Recovery Optimization: IBR and Hyaline-1S can go from N1 all the way back to N”.
 3.3 SCOT for Natarajan-Mittal Tree
@@
 Additionally, we need to check both tagged and flagged conditions for the current node, as (flagged) leaf nodes’s hazard pointer reservations need to be also verified.
-Due to complexity of the tree algorithm and very similar checks for optimistic traversals, we omit full pseudocode changes in this paper.
+Unlike Harris’ list, empirically, we found that recovery does not help improve performance for various key ranges, primarily because of the hierarchical structure of the tree (e.g., when there is a mismatch, likely, the tree has diverged substantially anyway).
+If ancestor no longer points to the original successor, we simply restart the search from the very beginning.
+Due to complexity of the algorithm and very similar checks for optimistic traversals, we omit pseudocode in this paper. In Section 5, we evaluate Natarajan-Mitttal Tree with SCOT.
 3.4 SCOT for Hash Map and Skip Lists
-SCOT for these data structures is largely identical to Section 4.2. Hash Maps are directly based on linked lists. Fraser’s skip list uses the same idea of logically deleted nodes as Harris’ list.
+SCOT for these data structures is largely similar to Section 3.2. Hash Maps are based on linked lists directly. Fraser’s skip list uses the same idea of logically deleted nodes as Harris’ list.
+4 Correctness
+We analyze the SCOT technique when applied to Harris’ linked list and the Natarajan-Mittal (NM) tree.
+Our goal is to establish two core properties under concurrent execution with SMR: (1) Safety: No thread accesses reclaimed or unlinked memory, and (2) Lock-Freedom: The system guarantees global lock-free progress - some thread completes its operation in a finite number of steps.
+We present invariant-based correctness arguments below.
+4.1 Model and Setup
+Our reasoning is based on a standard concurrent execution environment, assuming the following commonly accepted properties:
+  • Sequential consistency for all loads and stores.
+  • Atomic compare-and-swap (CAS) operations.
+  • Safe memory reclamation (SMR) with hazard pointers (HP):
+    • A node protected by a hazard pointer HP(T, i) is not reclaimed until all hazard reservations are cleared.
+    • Reclamation is conservative: memory is retired before reuse.
+Node Semantics:
+  • A node $n$ is logically deleted if isMarked(n.next) = true.
+  • A node is physically deleted if its predecessor no longer points to it.
+4.2 Definitions
+Let $H$ denote the Harris linked list under consideration, consisting of nodes $N_{1},N_{2},\dots$, each with a next pointer that may be marked to indicate logical deletion.
+  • Dangerous Zone $\mathcal{D}$: A maximal subsequence $\langle N_{k},\dots,N_{\ell}\rangle$ such that:
+    • $\forall i\in[k,\ell]$: isMarked($N_{i}$.next) = true,
+    • $\forall i\in[k,\ell-1]$: $N_{i}$.next = $N_{i+1}$,
+    • $N_{k-1}$.next = $N_{k}$.
+  • Validation Predicate: $\text{Valid}(a,b)\triangleq a.\texttt{next}=b$
+[…]
+4.3 Invariants
+Invariant 1 (Safe Traversal).
+A thread traversing a dangerous zone $\mathcal{D}$ must ensure at each step: $\text{Valid}(N_{i-1},N_{i})\wedge\texttt{HP(T, 3)}=N_{i}$
+Invariant 2 (ABA Prevention).
+A node protected by hazard pointers cannot be reclaimed or reused until all hazard reservations are cleared.
+Invariant 3 (Path Validation).
+Before accessing a node current in the NM tree, a thread must ensure that:
+  • the link ancestor $\rightarrow$ successor is valid, and
+  • current is reachable from successor.
+4.4 Structural Guarantees
+Lemma 1 (Ordered Node Removal).
+In Harris’ list and NM tree, nodes are always unlinked starting from the head of logically deleted chains. This prevents:
+  • Mid-chain removals that could violate traversal safety
+  • Reattachment of logically deleted nodes to other deleted nodes
+4.5 Traversal Strategy in Harris’ List
+SCOT maintains a bounded buffer of previously validated "safe" nodes during traversal. If validation fails, the thread:
+  1. 1.
+    Attempts to resume from the most recent valid node in buffer
+  2. 2.
+    Only restarts from head if all buffer entries are invalidated
+4.6 Memory Bounds
+Lemma 2 (Bounded Memory Overhead).
+The total memory overhead for hazard pointers is bounded and grows linearly with the number of threads.
+Proof.
+Let $N$ be the number of threads, $K$ the number of hazard pointers per thread, and $S$ the size of a node.
+Then the total hazard pointer memory usage is: $\text{Memory}=N\times K\times S$.
+In our SCOT-based implementation:
+  • The Harris list requires one more hazard pointer per thread than the Harris-Michael list, due to the need to traverse chains of logically deleted nodes safely.
+  • The Natarajan-Mittal tree operates within standard hazard pointer usage for tree traversal, requiring no additional hazard pointers beyond typical SMR-compatible designs.
+Since $K$ and $S$ are constants, the total memory overhead scales as $O(N)$, ensuring bounded and predictable memory usage. ∎
+4.7 Harris’ List Correctness
+Theorem 1 (HP Compatibility & Traversal Safety).
+Let thread $T$ traverse a dangerous zone $\mathcal{D}=\langle N_{k},\dots,N_{\ell}\rangle$ with:
+[…]
+[…]
+  • Validation: $\text{Valid}(N_{k-1},N_{k})$.
+Then $T$ never accesses a reclaimed node.
+Proof.
+  • Base case: $N_{k}$ is protected and validated via $N_{k-1}.\texttt{next}=N_{k}$.
+  • Inductive step: Before accessing $N_{i+1}$, $T$ validates $N_{i}.\texttt{next}=N_{i+1}$ and protects it.
+  • Fallback mechanism: If validation fails, the thread does not immediately restart from the head.
+Instead, it backtracks through a bounded buffer of previously validated safe nodes.
+It resumes traversal from the most recent entry in the buffer whose next pointer is still valid.
+Only if all buffered entries (buffer size is bounded) are invalidated (e.g., due to concurrent modifications), $T$ performs a full restart from the head.
+  • Termination: This buffered fallback approach preserves Invariant 1 and ensures that all accessed nodes remain protected and reachable, while reducing unnecessary restarts.
+Theorem 2 (ABA Prevention).
+During traversal, no node in $\mathcal{D}$ is reclaimed or reused.
+Proof.
+$N_{i}$ is protected via HP(T, 3). SMR guarantees prevent reclamation while protected. Invariant 2 holds. ∎
+Theorem 3 (Lock-Freedom of Harris’ List).
+SCOT-based Harris’ List is lock-free.
+Proof.
+Restarts occur only when validation fails, implying a successful CAS by another thread (e.g., physical unlink).
+Each such CAS represents system-wide progress, bounding the number of retries for any thread.
+Thus, at least one thread always makes progress in finite steps, satisfying lock-freedom.
+∎
+4.8 Correctness of NM Tree with SCOT
+Clarification of Roles:
+  • current: node being visited.
+  • successor: last untagged node on path to current.
+  • ancestor: lowest node that may unlink a flagged subtree.
+Theorem 4 (NM Tree Traversal Safety).
+Let thread $T$ traverse with:
+[…]
+[…]
+[…]
+[…]
+current is reachable from successor (ensured by hazard pointer protection of the path during traversal). Then current is safe to access.
+Proof.
+current is hazard-protected. Invariant 3 ensures the path to it is valid. A concurrent unlink would invalidate the link, forcing a restart. ∎
+Theorem 5 (Lock-Freedom of NM Tree).
+SCOT preserves NM Tree’s lock-freedom.
+Proof.
+Restarts occur only when:
+  • successor is tagged, or
+  • ancestor $\rightarrow$ successor becomes invalid (implying a concurrent structural modification).
+Each restart corresponds to a forward progress event (e.g., physical unlink via CAS). Contention is thus bounded, and lock-freedom is preserved. ∎
+4.9 Summary of Guarantees
+Type Description Applies To
+Invariants
+$\mathbf{I_{1}}$ (1) Safe traversal via validation Harris’ List, NM Tree
+$\mathbf{I_{2}}$ (2) ABA prevention Harris’ List, NM Tree
+$\mathbf{I_{3}}$ (3) Stable ancestor-successor path NM Tree
+Lemmas
+$\mathbf{L_{1}}$ (1) Ordered node removal Harris’ List, NM Tree
+$\mathbf{L_{2}}$ (2) Bounded memory overhead Harris’ List, NM Tree
+5 Evaluation
+Our benchmark reuses and substantially extends the test harness from [20], which already implements Harris-Michael List (HMList) and a few reclamation schemes: EBR, HP, and HE.
+We implemented Harris’ List (HList) and Natarajan-Mittal Tree (NMTree) [^5] We omit Ellen et al.
+Tree as it is slower and significantly diverges from Natarajan-Mittal Tree.
+We also omit skip lists and hash maps, as Harris’ vs.
+Harris-Michael lists are their fundamental building blocks and capture the differences already, e.g., Fraser’s vs.
+Herlihy-Shavit skip lists have exactly the same difference with respect to optimistic traversals.
+from scratch by carefully considering all well-known optimizations.
+We also added IBR [26] and Hyaline-1S [18] reclamation schemes into the benchmark.
+Finally, we implemented extra optimizations for HP, HE and IBR, which allow to greatly reduce overheads due to cache misses by capturing local snapshots of shared data in retire [18].
+(a) Key Range = 512
+(b) Key Range = 10,000
+Figure 7: Recovery optimization for SCOT: HList w/o restart recovery vs. HList w/ restart recovery (50% Read - 50% Write): higher is better.
+We did not implement HP++ [12] due to its substantial API differences (i.e., lack of an easy integration).
+HP++ is already known to perform mostly worse than HP for the same data structure [12], i.e., HP++ shows some performance benefits only when comparing different data structures (Harris-Michael HP list vs.
+Harris’ HP++ list).
+However, this comparison is largely irrelevant as we are now able to implement Harris’ list and other data structures in HP directly.
+However, where it is justified, we also indirectly compare our results to the results previously reported by HP++’s authors.
+We evaluate HMList, HList, and NMTree under different reclamation strategies: NR (no reclamation, leak memory), EBR, HP, IBR, HE, and HLN (Hyaline-1S).
+For lists, for readability, we use (New) to mark algorithms that use our SCOT method.
+For NMTree, everyone uses SCOT except NR/EBR.
+These algorithms were previously infeasible for HP/IBR/HE/Hyaline-1S or contained serious bugs in the available code, as discussed in [12, 3].
+We calibrated all memory reclamation schemes to provide the highest possible throughput while also minimizing the number of not-yet-reclaimed objects.
+We found that amortizing retire list scans with frequency of 128 works well for EBR, HP, HE, IBR and Hyaline-1S.
+Additionally, for EBR, IBR, HE and Hyaline-1S we select the epoch increment frequency which corresponds to the number of threads multiplied by 12.
+We run our experiments on Ubuntu 22.04.5 LTS using an AMD EPYC 9754 system with 128 physical cores (256 hardware threads w/ hyperthreading enabled), 384 GiB of RAM, and a maximum clock speed of 3.10 GHz.
+The benchmark is written in C++ and compiled using Clang++ 14.0.0 with -O3 optimizations because its compiled code tends to perform slightly better than the code generated by Ubuntu’s (default) GCC 11.4.0.
+For memory allocation, we use Microsoft’s mimalloc [15] since it scales much better in multi-threaded code compared to glibc’s stock malloc.
+(To stress test algorithms, we also successfully ran all tests with glibc’s stock malloc as it returns memory back to the OS more aggressively.)
+(a) Key Range = 16
+(b) Key Range = 512
+(c) Key Range = 10,000
+Figure 8: Linked List Throughput (50% Read - 50% Write): higher is better.
+(a) Key Range = 128
+(b) Key Range = 100,000
+Figure 9: NM Tree Throughput (50% Read - 50% Write): higher is better.
+We measure both throughput and memory overhead.
+Memory overhead is represented by the number of not-yet reclaimed objects, which are nodes that remain in an unreclaimed state due to the use of SMR.
+To calculate memory waste, we periodically record the average number of not-yet reclaimed nodes in every thread.
+We skip memory overheads for Hyaline-1S due to its global nature of reclamation, making it hard to calculate the number of unreclaimed objects locally per thread using this low-overhead mechanism which works for other SMR schemes.
+Each benchmark begins with prefilling the data structure with unique keys using 50% of the specified key range.
+We conduct 5 independent runs, each lasting 10 seconds.
+The median of all runs is used as the final measure for both throughput and memory overhead.
+(Across all experiments, the relative standard deviation remains mostly below 1%.) Our experiments use 1, 16, 32, 64, 128, 256, and 384 threads to evaluate performance for different levels of concurrency (384 threads represent oversubscription).
+Additionally, we test varying key ranges based on the data structure.
+For list, we use key ranges of 16, 512, and 10,000, while for the tree, we use key ranges of 128 and 100,000.
+In this section, we present results for a common workload, 50% read – 50% write (mixed reads and writes).
+In Appendix (Section 8), we additionally present results for 90% read – 10% write (read-dominated) and 50% insert – 50% delete (write-only) workloads.
+Both workloads exhibit largely similar trends.
+All throughput figures include a special NR (no reclamation) baseline which simply leaks memory and demonstrates the “upper bound” for performance.
+Note that in some cases, EBR or other schemes can still outperform this baseline, e.g., when the cost of reclaiming memory is much lower than that of performing fresh allocations.
+We also noticed that NR may outperform SMRs, including EBR, under oversubscribed scenarios.
+We first show how a recovery optimization for SCOT (Section 3.2.1) helps different schemes when using Harris’ list.
+In Figure 7, we compare HList with and without restart recovery – when the last safe node no longer points to the first unsafe node in the dangerous zone.
+We evaluate for two key ranges: 512 and 10,000, under four different reclamation schemes: HP, HE, IBR, and Hyaline-1S.
+(EBR does not use SCOT and thus the recovery optimization does not apply.) The recovery-enabled variant uses a bounded buffer (IBR, Hyaline-1S) to resume traversal from an intermediate safe node rather than restarting from the head on each retry, thereby avoiding repeated full-list traversals and improving throughput.
+For HP and HE, it is a one-shot recovery due to memory barrier overheads discussed in Section 3.2.1.
+Across all reclamation schemes, we observe consistent performance improvements with recovery at all thread counts.
+The benefits are especially pronounced for smaller key ranges (512), where contention is high, particularly at higher thread counts.
+Restart recovery yields up to 25% higher throughput compared to the variant without recovery.
+We then evaluate SCOT for different data structures and SMRs using a version with recovery.
+In Figure 8, we present throughput results for Harris-Michael and Harris’ list under different reclamation schemes in a workload consisting of 50% read and 50% write operations.
+We observe that Harris’ list consistently outperforms Harris-Michael for lower key ranges (16 and 512).
+As the key range increases to 10,000, the performance gap narrows but Harris’ list still maintains a throughput advantage.
+Generally, IBR and Hyaline-1S exhibit excellent performance on Harris’ list with SCOT traversal, matching or approaching that of EBR.
+HE and HP also show the relative benefits of SCOT traversal on Harris’ list, though their overall performance is somewhat lower compared to EBR, IBR, and Hyaline-1S, as expected.
+When comparing with previously reported HP++’s results, we observe that our approach with HP consistently outperforms HP++ across both small and large key ranges.
+Notably, for our highest key range (10,000), list implementations are twice as fast as that of HP++.
+In Figure 10, we measure the memory overhead by showing the number of retired but not-yet-reclaimed nodes for Harris-Michael and Harris’ lists.
+The HP implementation keeps the minimum amount of such objects.
+When comparing HP to HP++, we find that HP significantly reduces memory consumption (waste).
+Furthermore, HP consistently outperforms HP++ in terms of memory overhead when integrated with Harris’ and Harris-Michael List, making it a more efficient choice for memory reclamation.
+Figure 9 presents the throughput of the Natarajan-Mittal tree under a 50% read and 50% write workload.
+For the smaller key range (128), the baseline performance (without reclamation) is comparable to that of the various reclamation schemes, with only a small performance gap observed.
+In contrast, Figure 9(b) highlights the scalability of the tree at higher key ranges, where the Natarajan-Mittal tree achieves up to 240 million operations per second at 256 threads using the EBR scheme.
+Hyaline-1S also performs competitively, reaching 210 million operations per second, making it the closest in performance to EBR in this configuration.
+When comparing our results with HP++’s reported results, we observe that our implementation is roughly 4x faster.
+Moreover, for the key range of 100,000, the reported HP++ vs.
+EBR gap was around 30%, whereas our HP vs.
+EBR gap is less than 30% and our IBR/HE/HLN vs.
+EBR gap is less than 10%.
+(a) Key Range = 16
+(b) Key Range = 512
+(c) Key Range = 10,000
+Figure 10: Linked List Avg Not-Yet-Reclaimed Objects (50% Read - 50% Write): lower is better.
+(a) Key Range = 128
+(b) Key Range = 100,000
+Figure 11: NM Tree Avg Not-Yet-Reclaimed Objects (50% Read - 50% Write): lower is better.
+Figure 11 presents the memory overhead of the Natarajan-Mittal tree under a 50% read and 50% write workload.
+Across both Figure 11(a) and Figure 11(b), we observe consistent trends in memory usage.
+HP maintains a lower number of nodes in the limbo state, which we attribute to its strict and conservative reclamation guarantees.
+In contrast, EBR exhibits the highest memory overhead, reflecting its more relaxed and delayed reclamation behavior.
+Natarajan-Mittal tree with HP proves to be highly efficient in terms of memory consumption across all workloads.
+In contrast, HP++ exhibits nearly 5x higher memory consumption than HP at a 100,000 key range when integrated with the Natarajan-Mittal tree.
 6 Related Work
@@
 Timothy L.
-Harris [10] was the first to present a practical non-blocking lock-free linked list.
+Harris [10] was the first to present a practical lock-free linked list.
 Maurice Herlihy and Nir Shavit pointed out [11] that read-only optimistic search operations are feasible with Harris’ list.
@@
 Maged M.
-Michael proposed [13] to disable optimistic traversals by unlinking logically deleted nodes one by one, making it feasible to use the list with HP.
-Our work allows to support Harris’ original list with HP and take advantage of optimistic traversals.
+Michael proposed [13] disabling optimistic traversals by unlinking logically deleted nodes one by one, making it feasible to use HP.
+SCOT enables support for Harris’ original list with HP, HE, IBR, and Hyaline, while using optimistic traversals.
 A skip list is a data structure with an efficient search, insertion, and deletion operations, achieving (probabilistically) logarithmic time complexity.
-Lock-free versions of skip lists [11, 9] are very similar to Harris’ list and Harris-Michael list, except that they effectively consist of several sublists.
+Lock-free skip lists [11, 9] are very similar to Harris’ list and Harris-Michael list, except that they consist of several sublists.
 Consequently, challenges related to logical deletion and optimistic traversals are identical.
-Hash maps store and retrieve key-value pairs in constant time. A lock-free hash map [13] is, effectively, an array of linked lists, which can be based on either Harris’ or Harris-Michael approach.
+Hash maps store and retrieve key-value pairs in constant time. A lock-free hash map [13] is an array of linked lists, which can be based on either Harris’ or Harris-Michael approach.
 Binary search trees are also commonly used for efficient search, insertion, and deletion operations.
@@
 Timnat-Petrank’s original wait-free linked list, implemented in HP, is similar to Harris-Michael list.
-With our method, it can now leverage Harris’ optimistic traversals for faster performance.
+With SCOT, it can now use Harris’ optimistic traversals for faster performance.
 7 Conclusion
-We introduced safe concurrent optimistic traversals (SCOT), a new method which enables support for data structures that are normally incompatible with HP, IBR, HE, and other similar methods that lack the wide applicability support.
-Despite prior beliefs [12] of incompatibility of Natarajan-Mittal Tree and Harris’ List with HP, we have found that not only they are feasible with HP, IBR, and HE, but they also can largely retain performance benefits when comparing to equivalent data structures without optimistic traversals.
+We introduced safe concurrent optimistic traversals (SCOT), a new method which enables support for data structures that are incompatible with HP, IBR, HE, Hyaline-1S, and similar SMRs that lack the wide applicability support.
+Despite prior beliefs [12] of incompatibility of Natarajan-Mittal Tree and Harris’ List with HP, we have demonstrated that not only they are feasible with HP, IBR, HE, and Hyaline-1S, but they also can largely retain performance benefits when comparing to equivalent data structures without optimistic traversals.
 We hope that our work will help to spur further research and reevaluate the existing challenges related to safe memory reclamation because lack of the wide applicability property may be addressed by other means.